# Nonogram: Complexity of Inference and Phase Transition Behavior

### v1

@@ Abstract
Nonogram is a popular combinatorial puzzle (similar in nature to Sudoku or Minesweeper) in which a puzzle solver must determine if there exists a setting of the puzzle parameters that satisfy a given set of constraints. It has long been known that the problem of deciding if a solution exists is a computationally difficult problem. Despite this fact, humans still seem to enjoy playing it. This work aims to reconcile these seemingly contradictory facts by (1) analyzing the complexity of the inference problem for Nonogram (the problem of determining if there exists a puzzle parameter that can be inferred from the constraints without guessing) and (2) experimentally establishing the existence of a phase transition behavior for this inference problem. Our results show that the difficulty of the inference problem is largely determined by the density of filled cells (positive parameters) in a given puzzle. Along the way we implement an efficient encoding of a Nonogram board as a Boolean formula in Conjunctive Normal Form (CNF) through the use of regular expressions in order to make our experiments feasible.
@@ 1 Introduction
Nonogram is a combinatorial puzzle game played on a grid of cells that is initially empty. Each cell can be filled or empty in the solution. Each row and column is associated with a sequence integers specifying lengths of runs of filled cells in that row/column, termed a description. Players form a solution by assigning a subset of the cells to be filled such that every row and column description is satisfied. A unique aspect of Nonogram is that the solutions typically form a recognizable image (Figure 1).
(a) Initial Board
(b) Solved
Figure 1: Nonogram of Albert Einstein
@@
Due to its popularity, Nonogram is a well-studied puzzle. Prior work has been done to show the problem of consistency (the existence of a solution) NP-complete through parsimonious reduction[16] as well as with tools more specialized for combinatorial puzzles [6, 10, 12]. Various algorithmic approaches have been offered for solving Nonogram, including regular expression matching [2], evolutionary algorithms [1], and integer programming [8, 14] have all been used, along with a slew of approaches that work directly with the Nonogram puzzle (see https://webpbn.com/survey/ for an extensive review of solvers).
@@
As noted by Kaye while considering the complexity of the consistency problem for Minesweeper [7], obtaining a complexity result of NP-complete for consistency does not guarantee that there is not a polynomial-time approach to playing the game. Later work proposes the problem of inference – the existence of a cell that can have its value deduced by logical inference – as a more natural expression of solving a combinatorial puzzle. This problem has been shown co-NP-complete for Minesweeper [13].
@@
The study of phase transition behavior in combinatorial problems is useful for understanding the average-case complexity of problems that are infeasible for humans to solve in the worst case. Typically, a parameter of the combinatorial problem can be identified as a threshold, on each side of which a characteristic of the problem is dramatically different. The ratio of clauses to variables is such a parameter for Boolean satisfiability [9], and edge density a similar parameter for k-colorability of graphs [15]. This is a particularly interesting problem for combinatorial puzzles, as there is considerable interest in understanding what makes such puzzles hard so that their difficulty can be precisely tuned. For the puzzles of Sudoku and Minesweeper, phase transition behavior for the consistency problem has been shown [3, 11].
@@
In section 2, the problem of inference for Nonogram is defined and shown to be co-NP-complete. Given this complexity result, the popularity of Nonogram could be considered surprising. However, in section 3, phase transition behavior for the inference problem is demonstrated empirically, explaining this potential juxtaposition. In the process of demonstrating this phase transition, a reduction from Nonogram consistency to CNF formulae is developed, leveraging the expressibility of Nonogram descriptions as regular expressions. The formulas can be solved iteratively under different assumptions to solve the inference problem.
@@ 2 Inference is co-NP-complete
A consistent Nonogram board $N=(R,C,m,n)$ and mapping of cells to filled/empty/indeterminate $f$, where $R$ is the row descriptions and $C$ the column descriptions – each a sequence of integer sequences – and $m$ and $n$ the dimensions of the board.
@@
Does there exist a cell such that for all solutions to $N$, that cell is assigned to the same value of filled or empty (and is not already assigned to that value)?
For a cell to have its value inferred, it must get assigned to filled or empty. Further, the cell must take on this value for every possible solution. Otherwise, assignments to different cells may pertain to different solutions, leading to the value of one of those cells needing to be reassigned. Here we show that the INFERENCE problem for Nonogram is co-NP-complete. To this end, it is first shown that negative instances for INFERENCE can be verified in polynomial time. Next, the INFERENCE problem is proved co-NP-hard by a reduction from Boolean unsatisfiability.
@@
We first show INFERENCE $\in$ co-NP. For an $m\times n$ Nonogram board $N$, a certificate for a negative instance to the INFERENCE problem is the following: for each cell $c_{ij}$ two consistent fillings of $N$ are presented such that one has $c_{ij}$ filled and the other $c_{ij}$ empty. Thus, no cell can have its value inferred. A filled board can be checked for consistency by simply verifying that each description is satisfied. There are $2mn$ boards to check and $m+n$ descriptions for each board, thus the certificate can be verified in time polynomial in the size of the Nonogram board.
@@
$\bar{v}$$v$$\bar{v}^{\prime}$$v^{\prime}$$\bar{v}^{\prime\prime}$$v^{\prime\prime}$11111121121112113133131411111111111112112113113311211121112111111
@@ 3 Phase Transition
The reduction leverages the natural expressability of a Nonogram board as a regular expression. For an arbitrary description $[l_{1},l_{2},\dots,l_{t}]$ (assuming all $l_{i}>0$), the regular expression $0^{*}1^{l_{1}}0^{+}1^{l_{2}}0^{+}\dotsm 1^{l_{t}}0^{*}$ recognizes all strings satisfying the description. It follows that a deterministic finite automaton can be constructed recognizing the same language. Then, a Boolean formula is constructed describing the behavior of the automaton, such that the formula is satisfied if an only if the input string to the automoton is accepted and of the proper length (given by the size of the board). The encoding for a board is simply the conjunction of the Boolean formulas for the rows and columns. For a line of length $n$ with description $0^{*}1^{l_{1}}0^{+}1^{l_{2}}0^{+}\dotsm 1^{l_{t}}0^{*}$, there are $(5n+2)(t+1+\sum_{i=1}^{t}l_{i})-4$ clauses, $(14n+2)t+8n-2+(11n+2)\sum_{i=1}^{t}l_{i}$ total variables, and $(2n+1)(t+\sum_{i=1}^{t}l_{i})+n$ distinct variables [4]. Note that $t$, the number of description elements, and $\sum_{i=1}^{t}l_{i}$, the number of filled tiles, are both bounded by $n$. Thus, there are $O(n^{2})$ variables and $O(n^{2})$ clauses in the formula for an input description of a line with $n$ cells. The precise encoding algorithms developed can be found in [4].
Before solving, boards must be generated. As noted in literature, one must take care when defining a “random” instance of a problem [5]. For this analysis, only boards that are consistent are considered. To generate a random consistent board with filled cell density $\rho$ and dimension $N\times N$, a random number from zero to one is sampled uniformly for each cell. If the number is less than or equal to $\rho$, the cell is filled. Otherwise it is left as empty. Boards generated with filled cell density $\rho$ will have $\rho N^{2}$ filled cells in expectation. The descriptions associated with this filling can then be extracted to create a puzzle. As noted previously, this ensures that every board created has a solution. With a partially filled board generated, descriptions for each row and column can be created by finding the runs that occur in each of the rows and columns, in order. By convention, a column starts at the top and a row starts on the left (see column 3 of Fig. 3c). With the descriptions, the traditional Nonogram board has been created and is ready for encoding as a Boolean formula.
@@
On the horizontal axis is the filled cell density. For each density, the average proportion of filled cells inferred is plotted in black (the left vertical axis), with each point shape corresponding to a different size of puzzle. As expected, at lower densities hardly any of the filled cells can be inferred by logical inference. There appears to be a threshold between densities of 0.39 and 0.42 at which the average proportion of filled cells inferred quickly increases to almost one. This behavior becomes sharper as the size of puzzle increases.
It can be seen that for the 15x15 puzzles, the average proportion of filled cells actually decreases slightly from a non-zero amount at the beginning. This is likely the byproduct of the relatively small board size. With a filled cell density of 0.03, there are only 6.75 filled cells in expectation. As a result, on a particular board, it is possible to generate puzzles with even fewer cells, possibly creating entirely inferable puzzles. For larger dimension, the probability of generating such boards quickly drops, a pattern that can already be seen from the increase in dimension from 15x15 to 20x20 (which has 12 expected filled cells at a density of 0.03).
Also pictured is the computational effort required to solve the puzzles. Given that the underlying process is SAT solving, metrics relating to the SAT solving process are tracked. In this case, the average number of propagations by the solver for each board are measured. However, propagations is strongly associated with formula size, and the recurrences for formula size are parameterized by the size of the board. As such, the difficulty measure is plotted with empty point shapes and adjusted for board size, showing the percent of maximum average propagations for that board size. With this adjustment, the relationship between filled cell density and solving difficulty can be compared for different board sizes. Clearly, the hardness peaks around the phase transition threshold. This makes sense, as this is the region where the status as inferable or not inferable is most uncertain. Just as with the average proportion inferred, the spike in difficulty is starker for larger boards.
@@
The formula size and computational cost size clearly do not follow the same pattern. The number of clauses and variables is greatest above a filled cell density of 0.5, the puzzles for which solving is the easiest (fewest propagations).
@@ 4 Conclusion
This work build on complexity results fro the combinatorial puzzle Nonogram, considering problems directly related to playing the puzzle in a fashion that humans do. We have show that, assuming P $\neq$ NP, logical deduction to solve a Nonogram puzzle cannot be accomplished in polynomial time. To remedy this result with the popularity of the puzzle, we’ve demonstrated empirically phase transition behavior for the inference problem. In the process of doing so, we’ve demonstrated that the hardness of the puzzle peaks at the transition, and outside of the transition threshold, the computation required to solve puzzles is quite manageable.

### v2

@@ Abstract
Nonogram is a popular combinatorial puzzle similar in nature to Sudoku or Minesweeper in which a puzzle solver must determine if there exists a setting of the puzzle parameters that satisfy a given set of constraints. It has long been known that the problem of deciding if a solution exists is NP-complete. Despite this fact, humans still seem to enjoy playing it. This work aims to reconcile these seemingly contradictory facts by (1) analyzing the complexity of the inference problem for Nonogram (if there exists an unfilled cell whose value is forced) and (2) experimentally establishing the existence of a phase transition behavior for this inference problem. Our results show that the difficulty of the inference problem is largely determined by the density of filled cells (positive parameters) in a given puzzle. Along the way we implement an efficient encoding of a Nonogram board as a Boolean formula in Conjunctive Normal Form (CNF) through the use of regular expressions in order to make our experiments feasible.
@@ 1 Introduction
Nonogram is a combinatorial puzzle game played on an initially empty grid of cells, with some filled in for the solution. Each row and column is associated with a sequence of integers specifying lengths of runs of filled cells in that row/column, termed a description. Players form a solution by assigning a subset of the cells to be filled such that every row and column description is satisfied. A unique aspect of Nonogram is that the solutions typically form a recognizable image (Figure 1).
5112222221211222
(a) An empty Nonogram
5112222221211222
(b) A solved Nonogram
Figure 1: An example Nonogram of a smiley face.
@@
Due to its popularity, Nonogram is a well-studied puzzle. Prior work shows the problem of consistency (the existence of a solution) NP-complete through parsimonious reduction [16] as well as with tools more specialized for combinatorial puzzles [6, 10, 12]. Various algorithmic approaches exist for solving Nonogram, including regular expression matching [2], evolutionary algorithms [1], and integer programming [8, 14], along with a slew of heuristic approaches (see https://webpbn.com/survey/ for an extensive review of solvers).
@@
As noted by Kaye in the context of Minesweeper [7], obtaining an NP-complete result for consistency does not guarantee that there is not a polynomial-time approach to playing the game. Later work proposes the problem of inference – deciding the existence of a cell that can have its value deduced by logical inference – as a more natural expression of solving a combinatorial puzzle. After all, when players are handed a puzzle to play, they typically assume that there is a solution (consistency) and only one of them (uniqueness). The player is really interested in how difficult it is to make progress, and if any progress can be made at all without guessing, which is the inference problem. This problem has been shown co-NP-complete for Minesweeper [13]. Importantly, the inference problem is not the complement of the uniqueness problem which is known to be NP-complete [10]. An inference instance can be positive for puzzles with and without unique solutions (Figure 2). It is not hard to show that a puzzle has a unique solution if and only if every cell is inferable [4]. Furthermore, a satisfiable puzzle without any cells assigned may not have an inference possible.
11222211
(a) Inference possible with multiple solutions
122221
(b) Inference possible with a unique solution
Figure 2: Two puzzles in which inference is possible (the 2x2 square). In the first puzzle the pair of 1 clues creates multiple solutions, while the second puzzle only has one solution.
@@
The study of phase transition behavior in combinatorial problems is useful for understanding the average-case complexity of problems that are infeasible for humans to solve in the worst case. With this, a puzzle maker can tune their algorithm to generate puzzles that require more or less steps to solve. Typically, a parameter of the combinatorial problem can be identified as a threshold, on each side of which a characteristic of the problem is dramatically different. The ratio of clauses to variables is such a parameter for Boolean satisfiability [9], and edge density a similar parameter for k-colorability of graphs [15]. For the puzzles of Sudoku and Minesweeper, phase transition behavior for the consistency problem has been shown [3, 11].
@@
In section 2, the problem of inference for Nonogram is defined and shown to be co-NP-complete. In section 3, phase transition behavior for the inference problem is demonstrated empirically. In the process of demonstrating this phase transition, a reduction from Nonogram consistency to CNF formulae is developed, leveraging the expressibility of Nonogram descriptions as regular expressions. The formulas can be solved iteratively under different assumptions to solve the inference problem.
@@ 2 Inference is co-NP-complete
A consistent Nonogram board $N=(R,C,m,n)$ and mapping of cells to values $f$, where $R$ is the row descriptions and $C$ the column descriptions – each a sequence of integer sequences – and $m$ and $n$ the dimensions of the board.
@@
Does there exist a cell $c$ such that for all solutions $f$ to $N$, that $f(c)=v$, i.e., is $c$ mapped to the same value in every solution?
Here we show that the INFERENCE problem for Nonogram is co-NP-complete. To this end, it is first shown that negative instances for INFERENCE can be verified in polynomial time. Next, the INFERENCE problem is proved co-NP-hard by a reduction from Boolean unsatisfiability.
@@
We first show INFERENCE is a member of co-NP. For an $m\times n$ Nonogram board $N$, a certificate for a negative instance to the INFERENCE problem is the following: for each cell $c_{ij}$ two consistent fillings of $N$ are presented such that one has $c_{ij}$ filled and the other $c_{ij}$ empty. Thus, no cell can have its value inferred. A filled board can be checked for consistency by simply verifying that each description is satisfied. There are $2mn$ boards to check and $m+n$ descriptions for each board, thus the certificate can be verified in time polynomial in the size of the Nonogram board.
@@
$\bar{v}$$\bar{v}^{\prime}vv^{\prime}$$\bar{v}^{\prime\prime}$$v^{\prime\prime}$11111121121112113133131411111111111112112113113311211121112111111
@@ 3 Phase Transition
The reduction leverages the natural expressability of a Nonogram board as a regular expression. For an arbitrary description $l_{1},l_{2},\dots,l_{t}]$ (assuming all $l_{i}>0$), the regular expression $0^{*}1^{l_{1}}0^{+}1^{l_{2}}0^{+}\dotsm 1^{l_{t}}0^{*}$ recognizes all strings satisfying the description. It follows that a deterministic finite automaton can be constructed recognizing the same language. Then, a Boolean formula is constructed describing the behavior of the automaton, such that the formula is satisfied if an only if the input string to the automoton is accepted and of the proper length (given by the size of the board). The encoding for a board is a the conjunction of the Boolean formulas for the rows and columns. For a line of length $n$ with description $0^{*}1^{l_{1}}0^{+}1^{l_{2}}0^{+}\dotsm 1^{l_{t}}0^{*}$, there are $(5n+2)(t+1+\sum_{i=1}^{t}l_{i})-4$ clauses, $(14n+2)t+8n-2+(11n+2)\sum_{i=1}^{t}l_{i}$ total variables, and $(2n+1)(t+\sum_{i=1}^{t}l_{i})+n$ distinct variables [4]. Note that $t$, the number of description elements, and $\sum_{i=1}^{t}l_{i}$, the number of filled tiles, are both bounded by $n$. Thus, there are $O(n^{2})$ variables and $O(n^{2})$ clauses in the formula for an input description of a line with $n$ cells. The precise encoding algorithms developed can be found in [4].
Before solving, boards must be generated. As noted in literature, one must take care when defining a “random” instance of a problem [5]. For this analysis, only boards that are consistent are considered. To generate a random consistent board with filled cell density $\rho$ and dimension $N\times N$, a random number from zero to one is sampled uniformly for each cell. If the number is less than or equal to $\rho$, the cell is filled. Otherwise it is left as empty. Boards generated with filled cell density $\rho$ will have $\rho N^{2}$ filled cells in expectation. The descriptions associated with this filling can then be extracted to create a puzzle. As noted previously, this ensures that every board created has a solution. With a partially filled board generated, descriptions for each row and column can be created by finding the runs that occur in each of the rows and columns, in order. By convention, a column starts at the top and a row starts on the left (Fig. 4c). With the descriptions, the traditional Nonogram board has been created and is ready for encoding as a Boolean formula.
@@
On the horizontal axis is the filled cell density. For each density, the average proportion of filled cells inferred is plotted in black (the left vertical axis), with each point shape corresponding to a different size of puzzle. As expected, at lower densities few of the filled cells can be inferred by logical inference. There appears to be a threshold between densities of 0.39 and 0.42 at which the average proportion of filled cells inferred quickly increases to almost one. This behavior becomes sharper as the size of puzzle increases.
It can be seen that for the 15x15 puzzles, the average proportion of filled cells actually decreases slightly from a non-zero amount at the beginning. This is the byproduct of the relatively small board size. With a filled cell density of 0.03, there are only 6.75 filled cells in expectation. As a result, on a particular board, it is possible to generate puzzles with even fewer cells, possibly creating entirely inferable puzzles. For larger dimension, the probability of generating such boards quickly drops, a pattern that can already be seen from the increase in dimension from 15x15 to 20x20 (which has 12 expected filled cells at a density of 0.03).
Also pictured is the computational effort required to solve the puzzles. Given that the underlying process is SAT solving, metrics relating to the SAT solving process are tracked. In this case, the average number of propagations by the solver for each board are measured. However, propagation count is strongly associated with formula size, and the recurrences for formula size are parameterized by the size of the board. As such, the difficulty measure is plotted with empty point shapes and adjusted for board size, showing the proportion of average propagations scaled by the maximum result found for that board size. With this adjustment, the relationship between filled cell density and solving difficulty can be compared for different board sizes. Average propagation count appears to peak peaks around the phase transition threshold. This makes sense, as this is the region where the status as inferable or not inferable is most uncertain. Just as with the average proportion inferred, the spike in difficulty is starker for larger boards.
@@
Average clause and variable counts and computational cost appear to follow different patterns. The number of clauses and variables is greatest above a filled cell density of 0.5, the puzzles for which solving is the easiest (fewest propagations).
@@ 4 Conclusion
This work builds on complexity results for the combinatorial puzzle Nonogram, considering problems directly related to playing the puzzle in a fashion that humans do. We have shown that inferring Nonogram unfilled cell values is co-NP-complete. To remedy this result with the popularity of the puzzle, we have demonstrated empirically phase transition behavior for the inference problem. In the process we have demonstrated that the hardness of the puzzle peaks at the transition, and outside of the transition threshold, the computation required to solve puzzles is manageable.